OpenStudy (sleepyjess):

What is the rate of change between the interval of x = 0 and x = pi over two? graph below

OpenStudy (sleepyjess):

2 over pi pi over two pi over four four over pi

OpenStudy (johnweldon1993):

\[\large \frac{f(b) - f(a)}{b - a}\] So lets have a = 0 and b = pi/2 So when x = 0..what does y equal/ and when x = pi/2 what does y equal?

OpenStudy (sleepyjess):

0,3 and pi/2,5

OpenStudy (johnweldon1993):

Right...so \[\large \frac{f(b) - f(a)}{b - a} \rightarrow \frac{5 - 3}{\frac{\pi}{2} - 0} \rightarrow \frac{2}{\frac{\pi}{2}} \rightarrow 2 \times \frac{2}{\pi} \rightarrow \frac{4}{\pi}\]

OpenStudy (sleepyjess):

so it would be d?

OpenStudy (johnweldon1993):

D would be the correct answer...but do you see why?

OpenStudy (sleepyjess):

Yes I think I get it now. Thank you so much! :)